OpenStudy (anonymous):

integrate by parts: ⌡(x^3)/sqrt((x^2)+1))

OpenStudy (anonymous):

\(\huge \int \frac{x^3}{\sqrt{x^2+1}}dx \) let: \(\large u=x^3 \) and \(\large dv=\frac{1}{\sqrt{x^2+1}}dx \) so \(\large du=3x^2dx \) and \(\large v=arctanx \) so \(\large \int \frac{x^3}{\sqrt{x^2+1}}dx=v\cdot v - \int v du \) =\(\large (x^3)(arctanx) - \int arctanx \cdot 3x^2dx \) looks like you'll have to carry it out again... and again...

OpenStudy (anonymous):

ohhh ok. i see what i did wrong

OpenStudy (anonymous):

usually choose the u to be the one you take derivative so that the degree will be less after you take the derivative.

OpenStudy (anonymous):

yeah i got that part but i set dv\[=x(x ^{2}+1)^{\frac{ 1 }{ 2 }}\] and v=\[=(x ^{2}+1)^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

sorry..yes its to -1/2